OpenStudy (amanda0618):

Peter and his brother have a total of 12 fishing rods. Peter has 3 more than half of the amount his brother has. How many fishing rods does peters brother have? Thanks.

OpenStudy (superflytyguy):

1/2 of 12 is 6 right?

OpenStudy (alexandervonhumboldt2):

p+b=12 p=1/2b+3

OpenStudy (alexandervonhumboldt2):

1/2b+3=12-b

OpenStudy (superflytyguy):

Add 3 more

OpenStudy (alexandervonhumboldt2):

3/2b=9 b=6

OpenStudy (alexandervonhumboldt2):

brother has 6

OpenStudy (alexandervonhumboldt2):

so peter has alos 12-6=6

OpenStudy (superflytyguy):

@AlexandervonHumboldt2 is absolutely right